OpenStudy (anonymous):

A certain airline requires that the total outside dimensions (length + width + height) of a carry-on bag not exceed 34 inches. Suppose you want to carry on a bag whose length is twice its height. What is the largest volume bag of this shape that you can carry on a flight? (Round your answer to the nearest integer.)

OpenStudy (anonymous):

L + W + H = 34 3H + W = 34 W = 34-3H V = LHW = (2H)(H)(34-3H) = 68 H^2 -6 H^3 extreme value at dV/dH = 0 = 136 H - 18H^2 so divide by H or 18 H = 136 H= 7.6 W = 11.3 L = 15.2 L+W+H = 34.1 close enough, error due to rounding LWH = 1305 check that this is a maximum (d/dH)(dV/dH) = 136-36H = -138<0 yes and good night.